# EHC-MM: Embodied Holistic Control for Mobile Manipulation

### v1

@@ Abstract
Mobile manipulation typically entails the base for mobility, the arm for accurate manipulation, and the camera for perception. It is necessary to follow the principle of Distant Mobility, Close Grasping(DMCG) in holistic control. We propose Embodied Holistic Control for Mobile Manipulation(EHC-MM) with the embodied function of : By formulating the DMCG principle as a Quadratic Programming (QP) problem, dynamically balances the robot’s emphasis between movement and manipulation with the consideration of the robot’s state and environment. In addition, we propose the Monitor-Position-Based Servoing (MPBS) with , enabling the tracking of the target during the operation. This approach allows coordinated control between the robot’s base, arm, and camera. Through extensive simulations and real-world experiments, our approach significantly improves both the success rate and efficiency of mobile manipulation tasks, achieving a 95.6% success rate in the real-world scenario and a 52.8% increase in time efficiency.
@@ I INTRODUCTION
Mobile manipulation is expected to play a significant role in industrial and domestic. However, in many practical robot applications, the planning for the mobile base and robot arm is often separated. This separation in planning processes can lead to an unnatural approach overall, resulting in the loss of optimal solutions[1]. Consequently, this approach significantly increases the failure rate of planning and proves unsuitable for complex, unstructured environments. For mobile manipulation, Holistic Control, also known as Whole Body Control is essential and indispensable. Presently, in the context of mobile manipulation, two distinct challenges persist concerning the implementation of Holistic Control in physical robotic motions:
Lack of proprioception. Previous works mostly treat the entire robot as a single entity without considering the functional differences among its joints[1, 2, 3, 4]. While the robot is aware of the states of individual joints, it lacks the ability of self-awareness. Therefore, it’s crucial to coordinate control between the base and arm by considering the mobility and accuracy of each joint. The mobile base has greater mobility but lacks accuracy; the robot arm offers higher accuracy in manipulation but have limited workspace. That is why, for tasks involving movement and manipulation, we aim for the robot to the principle of Distant Mobility, Close Grasping (DMCG): prioritizing base movement when at a distance and focusing on arm grasping when close to the target [5], as shown in Fig. 1. Given inevitable perception and motion errors in real-world robots, integrating closed-loop systems with real-time planning and servoing is essential[5, 6].
@@
We designed to achieve embodied control. Specifically, accounts for both the reachability of the robot’s joints and the task objectives, allowing the robot to balance between movement and manipulation. Additionally, we developed a monitor-position-based servoing (MPBS) function to prevent the robot from losing track of the target during its operation.
@@ II-A Mobile Manipulation
Mobile manipulation is extensively explored across various domains in recent years. Notably, Chitta et al. [11] utilize visual, tactile, and proprioceptive sensor data to execute pick-and-place tasks with the PR2 robot. In addition, HERB [12] enables a robotic agent to identify grasping targets and navigate through complex environments. These contributions lay the foundation for mobile manipulation, emphasizing the segmentation of manipulation and navigation tasks. However, this partitioning in the planning domain is inherently artificial and introduces unnecessary complexity into the planning space.
@@ III ARCHITECTURE of EHC-MM
We present an architectural overview in Fig. 2. We divide the entire mobile manipulation into four parts: Real Robot, Target Perception, Motion Planning and Controller which are commonly observed in most robot frameworks. Our enhancements primarily focus within the blue rectangular box depicted in Fig. 2. The specific responsibilities of each component are elaborated below.
Embodied Weight Function: Enabling the robot to prioritize specific components of its motion at different distances.
Monitor and Position-based Servoing(MPBS) Controller: Allowing the robot to select the target end-effector pose at varying distances to avoid losing the target object.
GS-Net*: For generating a more stable grasping pose.
@@ III-A Kinematic Model and Control of the Mobile Manipulation
The resulting expression could also be written as the following differential kinematics,
@@
where and represent the coordinates of point and point , respectively. Evidently, inequality 3 is equivalent to
Building upon the model introduced by Jesse Haviland et al. [5, 13], we have enhanced the controller for mobile manipulation tasks by formulating it as a quadratic programming (QP) problem:
@@
is the decision variable, represents the difference between the desired and actual end-effector velocity – relaxing the trajectory constraint. incorporates joint velocity between the mobile base and the robot arm, incorporates manipulability maximisation and auxiliary performance tasks, incorporates an augmented manipulator Jacobian. represents the relative position of the obstacle point O to the virtual base. is the obstacle in the virtual base frame. is a gain term, and is a function which converts a homogeneous transformation matrix to a spatial twist.
III-B EHC Function
For mobile manipulation, the accuracy and motion abilities of each joint exhibit variations. Generally, robot arms exhibit higher accuracy but have limited workspace and higher computational complexity. Conversely, while the mobile base extends the operational workspace significantly, it generally introduces a greater degree of motion inaccuracy compared to robot arms. Hence, we designed the Embodied Holistic Control(EHC) Function, as shown below:
where
dynamically balances the emphasis between the movement and manipulation, allowing the robot to switch its focus between efficient movement and precise manipulation as required by the task at hand. In particular, represents the end-effector’s pose relative to the target from perception in Fig. 2. represents the set of end-effector poses with reachability greater than 95% for the given joint angles , while denotes the element within this set that is most closely aligned with the direction of the current task error , serving as a threshold for motion emphasis under the current task objective. In addition, represents the reachability of the robot arm’s end-effector. The activation function is incorporated into the robot control function, amplifying the emphasis on the specific motion. This enhancement improves the overall efficiency of motion planning. is the manipulability Jacobian.
By employing this approach, the robot can achieve the EHC: When the robot is positioned at a considerable distance from the target object, the mobile base primarily handles the task of approaching the target. As the robot nears the target and within an embodied threshold accounting for the robot’s joint states , the robot arm takes over as the primary actor in the manipulation process. Notably, this strategy enables the simultaneous and integrated planning of both locomotion and manipulation tasks, representing a significant advancement over traditional two-stage methods.
@@ III-C Monitoring-Position-Based Servoing
As illustrated in the Fig. 3(a), current Position-Based Servoing(PBS)[5] methods face the challenge of target loss during the reaching, which is particularly detrimental for reactive tasks. This is because, during the process of mobile manipulation, when the grasping orientation is not aligned with the forward-facing direction, such as when the orientation is downward or sideways, the end-effector of the robot arm changes its orientation prematurely during the reaching phase, resulting in a loss of visual contact with the target. Burgess et al.[21] propose improvements for grasping objects on the move. However, in their work, the object’s movement information is available to the robot. In real-world complex scenarios, the object’s positional information is only accessible to the robot through sensors. This necessitates that the robot should maintain the camera’s focus on the target during moving.
We utilize a controller that builds upon PBS and incorporates Monitoring-Position-Based Servoing (MPBS). This controller is designed to be parallel with the EHC function that dynamically balances the emphasis between monitoring and manipulation, based on . The MPBS is formulated as Equation 9, according to Rodrigues’ theorem.
where the is the end-effector target pose and is the pose for monitoring. represents the threshold value for switching between the grasping pose and the monitoring pose, as shown in Fig. 1. represents the rotation of axis towards the target, and represents the angle of rotation.
@@ IV-D Experiment-1: Randomly Reaching
Results and Analysis. We compared NEO(c), NEO(e), and EHC. In each set of experiments, the three approaches individually reached the same 50 random points. We conducted a total of 10 sets of experiments, thereby each method approaching 500 random points. The results of the experiment are displayed in the Table II. We find that our approach aligns with the DMCG principle, prioritizing base movement for navigation at a distance and arm manipulation for grasping or approaching tasks at close range. Additionally, due to the uncertainty associated with randomly generated points, some target poses may be challenging to reach. Hence, we set a maximum time limit of 30 seconds for each approach attempt. Otherwise, it is considered as a failure. Our method’s ability to coordinate movements and manipulation enables the robot to reach the desired target pose in a shorter time, resulting in fewer failures.
@@ IV-E Experiment-2: Sequentially Grasping Multiple Objects
Results and Analysis. We compare TSMM, NEO(c), NEO(e) and EHC. We conduct 15 sets of experiments, each grasping 45 objects. The results of the experiment are displayed in Table III, where NAN denotes that the value was not measured. Since, for the TSMM, measuring joint changes is meaningless as it includes two separate stages. Apart from the TSMM, we set a maximum grasping time of 30 seconds for each object. During real-world experiments, we find that the results in the real-world experiments align with those in simulations. EHC achieves shorter grasping times and higher success rates. we observe that the NEO(c) and NEO(e) still frequently perform major movements of the base even close to the target. Due to the limited precision of base movements, the robot faces difficulties in reaching the target smoothly, sometimes leading to failure. As for the TSMM, the robot frequently fails to grasp the most distant object because of the obstruction of the table. From Fig. 6, we can also visually observe that EHC enables smoother and faster mobile manipulation both in simulation and real-world experiments. Thus, EHC demonstrates excellent performance in real-world scenarios.
@@ V CONCLUSIONS
In this paper, we propose EHC-MM, leveraging the function of . By formulating the DMCG principle as a quadratic programming (QP) problem, we enable simultaneous planning of the robot’s joints, rather than relying on traditional two-stage mobile manipulation. The function dynamically balances the robot’s emphasis between movement and manipulation with the consideration of the robot’s state and environment, thereby improving both the success rate and efficiency of manipulation tasks. Additionally, is integrated into the MPBS, allowing the robot to maintain a balance between monitoring and manipulation, ensuring it does not lose track of the target during operation. Through simulation and real-world validation, we observe improved efficiency in mobile manipulation tasks using EHC. In real-world experiments, it achieves an impressive grasp success rate of 95.6%. We consider this performance to be highly commendable for real-world deployments. As robots become more complex with increasing degrees of freedom, EHC effectively coordinates joint actions within robot systems. In future work, we will further validate the efficiency of EHC across various robot platforms, including but not limited to mobile manipulation, dual-arm, and humanoid robots.

### v2

@@ Abstract
Mobile manipulation typically entails the base for mobility, the arm for accurate manipulation, and the camera for perception. The principle of Distant Mobility, Close Grasping(DMCG) is essential for holistic control. We propose Embodied Holistic Control for Mobile Manipulation(EHC-MM) with the embodied function of : By formulating the DMCG principle as a Quadratic Programming (QP) problem, dynamically balances the robot’s emphasis between movement and manipulation with the consideration of the robot’s state and environment. In addition, we propose the Monitor-Position-Based Servoing (MPBS) with , enabling the tracking of the target during the operation. This approach enables coordinated control among the robot’s base, arm, and camera, enhancing task efficiency. Through extensive simulations and real-world experiments, our approach significantly improves both the success rate and efficiency of mobile manipulation tasks, achieving a 95.6% success rate in real-world scenarios and a 52.8% increase in time efficiency.
@@ I INTRODUCTION
Mobile manipulation is expected to play a significant role in industrial and domestic applications. However, in many practical robot applications, the planning for the mobile base and robot arm is often separated. This separation in planning processes can lead to an unnatural approach, resulting in the loss of optimal solutions[1]. Consequently, this approach significantly increases the failure rate of planning and proves unsuitable for complex, unstructured environments. For mobile manipulation, Holistic Control, also known as Whole Body Control is essential and indispensable. Presently, in the context of mobile manipulation, two distinct challenges persist concerning the implementation of Holistic Control in physical robotic motions:
Lack of proprioception. Previous works mostly treat the entire robot as a single entity without considering the functional differences among its joints[1, 2, 3, 4]. Therefore, it’s crucial to coordinate control between the base and arm by considering the mobility and accuracy of each part. The mobile base has greater mobility but lacks accuracy; the robot arm offers higher accuracy in manipulation but has limited workspace. That is why, for tasks involving movement and manipulation, we aim for the robot to the Distant Mobility, Close Grasping (DMCG) principle: prioritizing base movement when at a distance and focusing on arm manipulation when close to the target [5], as shown in Fig. 1. Given inevitable perception and motion errors in the real world, integrating closed-loop systems with real-time planning and servoing is essential[5, 6].
@@
We design the to achieve embodied control. Specifically, accounts for both the reachability of the robot’s joints and the task objectives, allowing the robot to balance between movement and manipulation. Additionally, we developed a monitor-position-based servoing (MPBS) function to prevent the robot from losing track of the target during its operation.
@@ II-A Mobile Manipulation
Mobile manipulation is extensively explored across various domains in recent years. Notably, Chitta et al. [11] utilize visual, tactile, and proprioceptive sensor data to execute pick-and-place tasks with the PR2 robot. In addition, HERB [12] enables a robotic agent to identify grasping targets and navigate through complex environments. These contributions lay the foundation for mobile manipulation by emphasizing the segmentation of moving and manipulation tasks. However, this partitioning in the planning domain is inherently artificial and introduces unnecessary complexity into the planning space.
@@ III ARCHITECTURE of EHC-MM
We present an architectural overview in Fig. 2. We divide the entire mobile manipulation into four parts: Real Robot, Target Perception, Motion Planning and Controller which are commonly observed in most robot frameworks. Our enhancements primarily focus within the blue rectangular box depicted in Fig. 2.
@@ III-A Kinematic Model and Control of the Mobile Manipulation
The resulting expression could also be written as the following differential kinematics Equation 2, where represents the joint velocity vector of the robot.
@@
where and represent the coordinates of point and point , respectively. Evidently, inequality 6 is equivalent to
Building upon the model introduced by Jesse Haviland et al. [5, 13], we have further advanced the controller for mobile manipulation tasks by reformulating it as a quadratic programming (QP) problem, as described in Equation 8.
where is the decision variable, represents the difference between the desired and actual end-effector velocity, which relaxes the trajectory constraint. incorporates joint velocity between the mobile base and the robot arm, incorporates manipulability maximisation and auxiliary performance tasks, incorporates an augmented manipulator Jacobian.
The constraint function incorporates Equation 7 and joint velocity limits, and its upper bound is defined as:
where limits maximum and minimum values of the decision variable. The function describes the relative position relationship between a point in the robot link and the obstacle, as defined in Equation 11.
where represents the relative position of the obstacle point to the virtual base, computed as:
Here, is the obstacle in the virtual base frame. is a gain term, and is a function which converts a homogeneous transformation matrix to a spatial displacement.
III-B EHC Function
For mobile manipulation, the accuracy and motion abilities of each joint exhibit variations. Generally, robot arms exhibit higher accuracy but have limited workspace and higher computational complexity. While the mobile base extends the operational workspace significantly, it generally shows a greater degree of motion inaccuracy compared to robot arms. Hence, we designed the Embodied Holistic Control(EHC) Function, as shown below:
@@ III-B EHC Function
dynamically balances the emphasis between the movement and manipulation, allowing the robot to switch its focus between efficient movement and precise manipulation as required by target pose. In particular, represents the end-effector’s pose relative to the target from perception in Fig. 2. represents the set of end-effector poses with reachability greater than 95%, while denotes the threshold pose that is most closely aligned with the direction of the current target error , serving as a threshold for motion focus under the current task objective. In addition, represents the reachability of the robot arm’s end-effector. The activation function is incorporated into the robot control function, amplifying the emphasis on the specific motion. This enhancement improves the overall efficiency of motion planning. is the manipulability Jacobian.
By employing this approach, the robot can effectively adhere to the DMCG principle: When the robot is positioned at a considerable distance from the target object, the mobile base primarily handles the task of approaching the target, while the robot nears the target and within an embodied threshold accounting for the robot’s reachability , the robot arm takes over as the primary actor in the manipulation process. In particular, this strategy enables the simultaneous planning of both moving and manipulation tasks, representing a significant advancement over traditional two-stage methods.
@@ III-C Monitoring-Position-Based Servoing
As illustrated in the Fig. 3(a), current Position-Based Servoing(PBS)[5] methods face the challenge of target loss during the reaching process, which is particularly detrimental for reactive tasks. This is because, during the process of mobile manipulation, when the grasping orientation is not aligned with the forward-facing direction, such as when the orientation is downward or sideways, the end-effector of the robot arm changes its orientation prematurely during the reaching phase, resulting in a loss of visual contact with the target. Burgess et al.[21] propose improvements for grasping objects on the move. However, in their work, the object’s movement information is available to the robot. In real-world complex scenarios, the object’s positional information is only accessible to the robot through sensors. This necessitates that the robot should maintain the camera’s focus on the target during moving.
We utilize a controller that builds upon PBS and incorporates Monitoring-Position-Based Servoing (MPBS). This controller is designed to be parallel with the EHC function that dynamically balances the emphasis between monitoring and manipulation, based on . The MPBS is formulated as Equation 22,
where the is the end-effector target pose and is the pose for monitoring. represents the threshold value for switching between the grasping pose and the monitoring pose, as shown in Fig. 1.
@@ IV-D Experiment-1: Randomly Reaching
Results and Analysis. We compared NEO(c), NEO(e), and EHC. In each set of experiments, the three approaches individually reached the same 50 random points. We conducted a total of 10 sets of experiments, thereby each method approaching 500 random points. The results of the experiment are displayed in the Table II. We find that our approach aligns with the DMCG principle, prioritizing mobile base for moving at a distance, and robot arm for manipulation at close range. Additionally, due to the uncertainty associated with randomly generated points, some target poses may be challenging to reach. Hence, we set a maximum time limit of 30 seconds for each approach attempt. Otherwise, it is considered as a failure. Our method’s ability to coordinate movements and manipulation enables the robot to reach the desired target pose in a shorter time, resulting in fewer failures.
@@ IV-E Experiment-2: Sequentially Grasping Multiple Objects
Results and Analysis. We compare TSMM, NEO(c), NEO(e) and EHC. We conduct 15 sets of experiments, each grasping 45 objects. The results of the experiment are displayed in Table III, where NAN denotes that the value was not measured. Since, for the TSMM, measuring joint changes is meaningless as it includes two separate stages.
Apart from the TSMM, we set a maximum grasping time of 30 seconds for each object. During real-world experiments, we find that the results in the real-world experiments align with those in simulations. EHC achieves shorter grasping times and higher success rates. we observe that the NEO(c) and NEO(e) still frequently perform major movements of the base even close to the target. Due to the limited precision of base movements, the robot faces difficulties in reaching the target smoothly, sometimes leading to failure. As for the TSMM, the robot frequently fails to grasp the most distant object because of the obstruction of the table. From Fig. 6, we can also visually observe that EHC enables smoother and faster mobile manipulation both in simulation and real-world experiments. Thus, EHC demonstrates excellent performance in real-world scenarios.
@@ V CONCLUSIONS
In this paper, we propose EHC-MM, leveraging the function of . By formulating the DMCG principle as a quadratic programming (QP) problem, we enable simultaneous planning of the robot’s joints, rather than relying on traditional two-stage mobile manipulation. The function dynamically balances the robot’s emphasis between movement and manipulation with the consideration of the robot’s state and environment, thereby improving both the success rate and efficiency of manipulation tasks. Additionally, is integrated into the MPBS, allowing the robot to maintain a balance between monitoring and manipulation, ensuring it does not lose track of the target during operation. Through simulation and real-world validation, we observe improved efficiency in mobile manipulation tasks using EHC. In real-world experiments, it achieves an impressive grasp success rate of 95.6%. The results demonstrate that the proposed method is highly effective for real-world deployments.
VI ACKNOWLEDGMENT
The work is supported by the National Natural Science Foundation of China (No. 62176004), Beijing Natural Science Foundation(No. L222012), Wuhan East Lake High-Tech Development Zone(also known as the Optics Valley of China, or OVC) National Comprehensive Experimental Base for Governance of Intelligent Society, Intelligent Robotics and Autonomous Vehicle Lab (RAV), Beijing Samsung Telecom R&D Center.